The bases of a softball diamond are 60 feet apart. How far is it from home plate to second base?
Answer by wilft1(217)

use pathagorias theorm for this a^2+b^2=c^2
60 x 60 is 3600
so
3600 + 3600 = 7200
the square root of 7200 is 84.85
so the distance between home plate and 2nd base is 84.85 feet